OpenStudy (anonymous):

*** help*** a new exhibit is scheduled to open at the local museum. museum officials expect that 5000 people will visit the exhibit in its first week, and that the number of visitors will drop by 10 people per week after the first week during the first 6 months. find the total number of visitors expected in the exhibits first 7 weeks

OpenStudy (mathstudent55):

In the first week there will be 5000 visitors. In the second week, there will be 10% fewer visitors than in the first week. Can you calculate how many visitors there will be in the second week?

OpenStudy (anonymous):

5000/.10=50000

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

No. In math, in addition to solving a problem using some method, you need to look at your answer and see if it makes sense. If there will be 10% FEWER visitors in the second week, how can you end up with 10 times more visitors?

OpenStudy (mathstudent55):

This is how you do it.

OpenStudy (mathstudent55):

In the first week, there are 5000 visitors. Call the initial number, 5000, 100%. In the second week, there will be 10% fewer visitors. That means in the second week, there will be 100% - 10% = 90% of the 5000 visitors.

OpenStudy (mathstudent55):

Now we calculate 90% of 5000. 90% of 5000 = 90/100 * 5000 = 0.90 * 5000 = 4500

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

Ok so far?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

but the number that you're getting is decreasing. the answer choices are, 34,730, 24,850, 34790, 29,850

OpenStudy (mathstudent55):